tahmina:

susan begins counting backward from 1298 by 4's saying one number every 5 sec. At the same time Jim begins counting forward from 171 by 3's saying one number every 5 sec. what number they will both sat at the same time.

celticcat:

THese are 2 arithmetic sequences , so:- nth term = 1298 - 4(n-1) - for susan's sequence nth term = 171+ 3 (n-1) - for jim's sequencq when terms are equal:- 1298 - 4(n -1) = = 171 + 3(n - 1) 1298 -4n + 4 = 171 + 3n - 3 7n = 1298 + 4 - 171 + 3 7n = 1134 n = 162 so the same number is said on the 162nd term which is 1298 - 4(162-1) = 654 check this for the othere squence 171 +3(162-1) = 654 also So the answer is 654